OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! (08.03)Solve the system of equations and choose the correct answer from the list of options. x − y = 7 y = 3x + 12 2 over 19 comma 2 over 33 negative 2 over 19 comma negative 33 over 2 negative 19 over 2 comma negative 33 over 2 19 over 2 comma 33 over 2

OpenStudy (tommynaut):

Do you understand what the question is asking for? When you solve 2 equations at the same time, you're solving for 2 different variables (letters) at the same time; this means you're trying to find the value of x and the value of y that makes each equation true. For example, the equation 2x = 4 can be solved with the solution that x = 2. This is only one equation, however. When we solve for 2 equations simultaneously (at the same time) we need to consider just one equation first. There are a few ways of solving simultaneous equations - the easiest way in your situation is to use SUBSTITUTION. Let's work through an example. x + y = 5 ... (equation 1) x - y = 1 ... (equation 2) First of all, we choose an equation (it doesn't matter which!). Let's choose equation (2). Let's choose a variable (again, it doesn't matter which one we choose!) that we want all by itself... I'm going to choose x. If we rearrange the equation, we get x = 1 + y (we just added y to both sides of the equation) We then SUBSTITUTE this "new" equation into (1), so we get (1+y) + y = 5, which we can then solve 1 + 2y = 5 2y = 4 y = 2 Now that we have y = 2, we can SUBSTITUTE that back into either equation to find x. Let's choose (2) again (it doesn't matter which one we choose here either!) x - y = 1, so x - 2 = 1 x = 3 So we end up with x = 3, y = 2. Try a similar approach in your question.